OpenStudy (anonymous):

Find the parameterization of the semi-circle given by x^2+y^2=255, y<=0 in the xy-plane starting at the point (-15, 0) and ending at (15, 0)?

OpenStudy (anonymous):

Is it 225? (rather than 255) Just the usual cos t sin t and restrict the domain between 0 and pi.

OpenStudy (anonymous):

\[x=a \cos(t) ............y=bsin(t)......... using 0<t<2\pi\]

OpenStudy (anonymous):

\[x=\sqrt{255}\cos(t)...............y=\sqrt{255}\sin(t)\]

OpenStudy (anonymous):

i believe estudier is correct if y^2+x^2=225 cos^2(t)+sin^2(t)=1 15x=cos(t) 15y=sin(t) Then you rescrict the domain to tE[0,pi]

OpenStudy (anonymous):

Sorry it's 225 not 255

OpenStudy (anonymous):

so x=15cos(t) and y=15sin(t) from 0 to pi

OpenStudy (anonymous):

when you said y<=0 the semi circle is below the x axis this means at a certain point y needs to be -15 so we restrict the domain from t E [pi,2pi] not [0,pi]

OpenStudy (anonymous):

I see. I understand now. Thanks!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Ah, lower not upper half, good spot...